OpenStudy (anonymous):

A stone is thrown vertically downward from a cliff 186.0-m tall. During the last half second of its flight, the stone travels a distance of 46.9 m. Find the initial speed of the stone. Note: I won't be able to respond for a few hours

OpenStudy (anonymous):

Is this just basic distance time speed formula? If so...piece of cake. :D

OpenStudy (anonymous):

Nm

OpenStudy (anonymous):

They don't give a time

OpenStudy (anonymous):

crapola

OpenStudy (anonymous):

back to the drawing board then

OpenStudy (anonymous):

figured it out I think

OpenStudy (anonymous):

divide 186 by 46.9

OpenStudy (anonymous):

multiply the result by .5

OpenStudy (anonymous):

1.98

OpenStudy (anonymous):

1.98 in minutes and seconds is...

OpenStudy (anonymous):

1 minute and 58 seconds

OpenStudy (anonymous):

I THINK I'm no physicist...

OpenStudy (anonymous):

Speaking of which, I must admit @Abmon98, I MIGHT be stalking you for help in Chemistry

OpenStudy (anonymous):

wait, wait, wait I read it as half a minute, not half a second crapola

OpenStudy (abmon98):

haahah sure i will help you give me 5 more minutes

OpenStudy (anonymous):

well...that would just mean... 1 second and 3 half seconds and thanks @Abmon98

OpenStudy (abmon98):

s=46.9 u=xm/s a=10m/s^2 t=0.5 second ut+1/2at^2=s that initial velocity is the final velocity when the stone was travelling ((186-46.9)=139.1m at s=139.1 m a=10m/s^2 v=91.300 m/s use Vt-1/2at^2=s solve for t and then use u+v/2)*t=s